OpenStudy (anonymous):

v =7i -2j, w= -3 - j *** find v.w and v.v

OpenStudy (anonymous):

Dot product? If it's dot product: v . w = -21i +2j v . v = 49i +4j

OpenStudy (anonymous):

i think its multiply v X w v X v

OpenStudy (gw2011):

(v)(w)=(7i-2j)(-3-j) =-21i-7ij+6j+2j^2 =-7i(3+j)+2j(3+j) =(-7i+2j)(3+j) (v)(v)=(7i-2j)(7i-2j)=49i^2-28ij+4j^2